# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

### v1

@@ 3.1 Overall Framework
The architecture of Kairos, shown in Figure 2, consists of three main components. First, the input time series is tokenized into a sequence of embeddings by our Mixture-of-Size Dynamic Patching (MoS-DP) module to extract multi-granularity local information. These embeddings are then processed by a Transformer encoder, which uses our proposed Instance-Adaptive Rotary Position Embedding (IARoPE) to model complex temporal relationships. Finally, a Transformer decoder utilizes a multi-patch prediction strategy for forecasting. The details of each key component, including MoS-DP, IARoPE, and the multi-patch prediction strategy, are presented in the following sections.
3.2 Mixture-of-Size Dynamic Patching (MoS-DP)
In contrast to time-series foundation models such as Time-MoE (Shi et al., 2024) and Moirai-MoE (Liu et al., 2024b), which directly replace the original feed-forward network (FFN) layers in Transformer blocks with Mixture-of-Experts (MoE) modules (Jacobs et al., 1991; Shazeer et al., 2017), Kairos takes inspiration from MoE to tackle the intrinsic heterogeneity of time series. To improve local information extraction, we introduce the MoS-DP module, which performs routing at the level of each time series patch and thereby effectively addresses the challenge of adapting foundation models to diverse datasets with varying information granularities. MoS-DP unfolds in three stages: Patchify at Coarsest Granularity, Dynamic Patch Router, and Dynamic Patch Fusion.
Patchify at Coarsest Granularity. In this stage, the entire time series is partitioned based on the maximum patch size. We begin with an initial partition of the sequence. A set of $S$ patch sizes is defined in advance, denoted as $\{p_{1},\ldots,p_{S}\}\in\mathbb{R}^{S}$, where $p_{1}<p_{2}<\ldots<p_{S}$. We first tokenize the input time series $\bm{x}_{1:T}\in\mathbb{R}^{T}$ using non-overlapping patching using the coarsest patch size $p_{S}$. The series is divided into $N=\lceil\frac{T}{p_{S}}\rceil$ coarsest patches. If the original series length $T$ is not an exact multiple of $p_{S}$, $\bm{x}_{1:T}$ is augmented with left-padding of zeros as needed. This process yields a sequence of $N$ coarsest patches, $\{\bm{x}_{n}^{p_{S}}\}_{n=1}^{N}$, where each patch has a uniform length of $p_{S}$.
Dynamic Patch Router. To perform the final multi-granularity information fusion, we select a set of appropriate granularities for each coarsest patch $\bm{x}_{n}^{p_{S}}$. Inspired by the MoE paradigm, we employ multiple FFNs as experts, with each expert dedicated to extracting features at a unique granularity level from the set $\{p_{1},\ldots,p_{S}\}$. Furthermore, we introduce $Z$ null experts, which perform no computation when selected, allowing the number of activated experts to be dynamic for each input, facilitating more flexible modeling of heterogeneous time series. The routing process is formulated as
$g_{n,i}=\begin{cases}s^{\prime}_{n,i},&s^{\prime}_{n,i}\in\mathrm{TopK}(\{s^{\prime}_{n,j}|1\leqslant j\leqslant S+Z\},K)\ \text{and}\ 1\leqslant i\leqslant S\\ 0,&\text{otherwise.}\end{cases}$ (1)
$s^{\prime}_{n,i}=\frac{\exp(s_{n,i}+\bm{b}_{i})}{\sum_{j=1}^{S}\exp(s_{n,j}+\bm{b}_{j})},\quad\text{where }s_{n,i}=\mathbf{W}_{i}\bm{x}_{n}^{p_{S}}.$ (2)
Here $g_{n,i}$ represents the gating value for the $i$-th expert with respect to the $n$-th coarsest patch; $s_{n,i}$ is the affinity score between patch and the expert, and $\mathbf{W}_{i}$ is the trainable weight matrix for the $i$-th expert. The bias term $\bm{b}_{i}$ is utilized for an Auxiliary-Loss-Free Load Balancing method (Liu et al., 2024a) and is discussed in Appendix C.3.1.
The coarsest patch $\bm{x}_{n}^{p_{S}}$ will then be partitioned into smaller finest patches. The size of these finest patches, $p_{k}$ is determined by the finest granularity selected by the router for this specific coarsest patch, i.e., $p_{k}=\min\{p_{i}|g_{n,i}>0\}$. Note that this finest granularity can differ for each coarsest patch $\bm{x}_{n}^{p_{S}}$. All $N$ coarsest patches are re-segmented in parallel based on their respective routing decisions. Compared to uniformly segmenting the entire sequence with a fixed size (Chen et al., 2024; Zhang et al., 2024) as discussed in Section 2, our method offers greater flexibility to adapt to temporal sequences with varying information densities.
Dynamic Patch Fusion. In the final stage, we yield the embedding for each finest patch by fusing information from multiple granularities. This process ensures that the final token representations are aware of the context at various temporal scales. Specifically, for the $n$-th coarsest patch $\bm{x}_{n}^{p_{S}}$, which was determined in the previous stage, it is first partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches. Here, $p_{k}$ is the finest patch size selected by the Dynamic Patch Router for this coarsest patch. We denote the $m$-th finest patch within the $n$-th coarsest patch as $\bm{x}_{n,m}^{p_{k}}$. Each such finest patch corresponds to a single token for the subsequent Transformer encoder. The embedding $e_{n,m}$ for the finest patch $\bm{x}_{n,m}^{p_{k}}$ is calculated as the weighted summation of semantic information originating from all activated granularities, formulated as:
$\bm{e}_{n,m}=\sum_{i=k}^{S}\alpha_{n,i}\cdot\text{MLP}_{i}(\mathcal{A}(x_{n,m}^{p_{k}},p_{i})),\ \text{where}\ \ \alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=k}^{S}g_{n,j}},\ \ \bm{e}_{n,m}\in\mathbb{R}^{D_{h}},$ (3)
where $\mathcal{A}(x_{n,m}^{p_{k}},p_{i})$ identifies the ancestor patch of size $p_{i}$ that contains the finest patch $x_{n,m}^{p_{k}}$, as illustrated in Figure 2. We provide a formal definition of $\mathcal{A}(\cdot,\cdot)$ and a detailed example in Appendix C.3.2. $\text{MLP}_{i}$ denotes the expert that extracts the embedding for patch size $p_{i}$. The term $\alpha_{n,i}$ represents the normalized gating value from the router for the $n$-th coarsest patch, which dynamically weighs the importance of each granularity. $D_{h}$ is the hidden dimension. Following Eq. 3, $M$ embeddings are extracted from the coarsest patch $\bm{x}_{n}^{p_{S}}$ via the Dynamic Patch Fusion (DPF) module as:
$\text{DPF}(\bm{x}_{n}^{p_{S}},\bm{g}_{n})=[\bm{e}_{n,1},\bm{e}_{n,2},\ldots,\bm{e}_{n,M_{n}}]\in\mathbb{R}^{M_{n}\times D_{h}},\quad\bm{g}_{n}=(g_{n,1},g_{n,2}\ldots,g_{n,S}).$ (4)
Eventually, the original time series $\bm{x}_{1:T}$ is transformed into a sequence of embeddings by concatenating all DPF module outputs as $\mathbf{E}=\text{DPF}(\bm{x}^{p_{S}}_{1},\bm{g}_{1})||\text{DPF}(\bm{x}^{p_{S}}_{2},\bm{g}_{2})||\ldots||\text{DPF}(\bm{x}^{p_{S}}_{N},\bm{g}_{N})$ with sequence length $T^{\prime}=\sum_{n=1}^{N}M_{n}$, where $M_{n}$ represents the number of finest patches that the $n$-th patch in $\bm{x}^{p_{S}}_{1:N}$ is segmented into by the Dynamic Patch Router. This final sequence $\mathbf{E}\in\mathbb{R}^{T^{\prime}\times D_{h}}$ would serve as the input to the subsequent Transformer module of Kairos. This design allows the model to use small patches for in-depth analysis where local details are rich and large patches for efficient modeling where information is stable, as demonstrated in Section 4.4.1.
3.3 Instance-Adaptive Rotary Position Embedding (IARoPE)
Rotary Position Embedding (RoPE) (Su et al., 2024) encodes relative positions by rotating hidden vectors within Transformer layer, such as a query ($\bm{q}$) or a key ($\bm{k}$), which we denote generically as $\bm{z}$. The operation is formulated as $f_{\text{RoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta_{\text{init},j}}$, where the base RoPE frequencies $\theta_{\text{init},j}=b^{-2j/D_{h}}$ are determined by a fixed hyperparameter $b$. A detailed derivation of the RoPE calculation is provided in Appendix C.4. However, these fixed frequencies impose a single, uniform temporal scale across all data. This rigidity prevents the model from effectively modeling temporal relationships for series with diverse intrinsic structures, such as different seasonalities or trends. To address this limitation, we propose Instance-Adaptive Rotary Position Embedding (IARoPE), which dynamically tailors customized positional encodings for each time series instance.
The core idea of IARoPE is to modulate the base RoPE frequencies $\theta_{\text{init}}$ using spectral features extracted from each instance. Specifically, Given an input time series instance $\mathbf{x}_{1:T}$, we first extract its frequency features by applying the Fast Fourier Transform (FFT) and select the first $\omega$ low-frequency components (their magnitudes) to form a feature vector $\mathbf{x}_{\text{fft}}$, where $\omega$ is a hyperparameter. We focus on low frequencies as they often capture dominant trends and periodic patterns crucial for understanding the temporal structure of instance (Zhou et al., 2022). Then we feed $\mathbf{x}_{\text{fft}}$ into a small MLP to predict instance-specific modulation parameters: a scaling factor $\mathbf{\gamma}\in\mathbb{R}^{D_{h}/2}$ and a bias factor $\mathbf{\beta}\in\mathbb{R}^{D_{h}/2}$. These parameters then adapt the base RoPE frequencies $\theta_{\text{init}}\in\mathbb{R}^{D_{h}/2}$ via an element-wise affine transformation to produce the adaptive frequencies $\theta^{\prime}\in\mathbb{R}^{D_{h}/2}$:
$\theta^{\prime}_{j}=\gamma_{j}\odot\theta_{\text{init},j}+\beta_{j}\quad\text{for }j=0,1,\ldots,D_{h}/2-1,$ (5)
where $\odot$ denotes the Hadamard product. The resulting $\theta^{\prime}$ incorporates learned information from the instance’s frequency characteristics. Finally, these adaptive angles $\theta^{\prime}$ replace the fixed $\theta$ in the RoPE calculation for the current instance:
$f_{\text{IARoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta^{\prime}_{j}},$ (6)
This mechanism adapts the positional encoding to the frequency profile of each time series instance, potentially offering a more accurate representation of relative temporal positions. We demonstrate the effectiveness of this adaptation in Section 4.4.2.
3.4 Multi-Patch Prediction
Kairos employs multi-patch prediction to effectively mitigate the cumulative errors inherent in autoregressive processes. Specifically, after the encoder outputs the hidden representation $\bm{h}^{\text{en}}\in\mathbb{R}^{T^{\prime}\times D_{h}}$, a Transformer decoder is employed to forecast future values. We introduce a set of $J$ learnable forecast tokens, denoted as $Q_{\text{forecast}}=\{o_{1},\dots,o_{J}\}$. These tokens are passed as input to decoder, which processes them simultaneously to produce a sequence of output embeddings $\bm{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$:
$\bm{h}^{\text{de}}=\text{Decoder}(Q_{\text{forecast}},\bm{h}^{\text{en}}).$ (7)
@@ 3.4 Multi-Patch Prediction
Each output embedding $\bm{h}^{\text{de}}_{j}$ is then passed through a shared prediction head (a residual FFN block (Das et al., 2024)) to forecast a patch of the future of length $H_{s}=H/J$ as
$\hat{\bm{x}}_{T+jH_{s}+1:T+(j+1)H_{s}}=\text{ResidualFFN}(\bm{h}^{\text{de}}_{j}),\quad j=\{1,\dots,J\}$ (8)
This multi-token design is motivated by the limitations of two common forecasting paradigms. First, compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches at once enhances computational efficiency and mitigates cumulative error by reducing iterative steps. Second, unlike models that forecast a larger patch (Das et al., 2024), our approach provides greater flexibility for shorter prediction horizons. It can construct a precise forecast by aggregating outputs from the necessary number of tokens, avoiding the potentially imprecise truncation of a larger-than-needed output. We provide a detailed analysis of this design in Appendix A.

### v2

@@ 3.1 Overall Framework
The architecture of Kairos, shown in Figure 2, consists of three components. First, the input time series is tokenized by the Mixture-of-Size Encoder to extract multi-granularity local representations via dynamic granularity selection. These embeddings are then processed by a Heterogeneity-Aware Transformer equipped with Dynamic Rotary Position Embedding (DRoPE), a granularity-aware positional encoding that performs instance-wise modulation and adapts to dynamic patching tokenization. Finally, the Multi-Patch Decoder generates forecasts by predicting multiple future patches in parallel, mitigating cumulative errors. The details of these components are presented in the following sections.
3.2 Mixture-of-Size Encoder
Prior to the Heterogeneity-Aware Transformer (Section 3.3), the Mixture-of-Size Encoder adaptively tokenizes the raw time series $\mathbf{X}$ through a three-stage process. First, the series is partitioned into fixed-size segments, where a router dynamically selects optimal temporal granularities for each segment. Next, for each activated granularity, the segment is simultaneously patchified and projected into the latent space by the expert network corresponding to that granularity. Finally, these multi-granularity embeddings are fused into a unified token sequence representing the input series.
Notably, the Mixture-of-Size Encoder is both sparse and dynamic. A gating mechanism equipped with null experts (described below) adaptively selects relevant granularities for each segment, enabling sparse expert activation and segment-specific granularity adaptation.
Coarse Segmentation. As shown in Figure 1(b) in Section 1 where time series exhibits inherent temporal heterogeneity, we first apply coarse segmentation to the input time series $\mathbf{X}$ and process each segment independently in subsequent tokenization and encoding stages. Specifically, $\mathbf{X}$ is partitioned into a sequence of non-overlapping coarse segments $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{N}]$, where each segment $\mathbf{x}$ has a fixed length $P$ and the total number of segments is $N=\lceil{T}/{P}\rceil$. In the following, we operate within each segment $\mathbf{x}$ and therefore omit the segment index for notational simplicity.
Granularity. For a given time series segment $\mathbf{x}$, we adopt a multi-granularity modeling perspective. Granularity refers to the temporal resolution at which the segment is tokenized and encoded, as defined below.
Definition 3.1 (Granularity Level).
A granularity level $i\in\{1,\ldots,G\}$ is defined by a specific patch size $P_{i}$ and its corresponding encoding network $\pi_{i}$, under which the time series segment $\mathbf{x}$ is patchified and mapped into a latent representation.
Intuitively, finer granularity levels capture high-frequency local variations, while coarser granularity levels emphasize low-frequency global trends. At each granularity level $i$, the patchification and embedding operations described below are applied independently and in parallel, producing representations that characterize the segment at the corresponding temporal resolution.
Step 1: Top-$K$ Granularity Routing. Applying encoding at an appropriate granularity level for each time series segment is non-trivial due to the inherent temporal heterogeneity of time series data. To address this challenge, we introduce a dynamic routing mechanism equipped with a lightweight router that adaptively selects an optimal subset of granularity levels for subsequent encoding of a given segment. Inspired by the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991; Shazeer et al., 2017), we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ null experts that perform no computation. The inclusion of null experts allows the model to implicitly skip encoding at certain granularities, enabling adaptive control over the effective number of active granularities for each input segment. The routing weights are computed via a learnable scoring function. Specifically, for each expert $i\in\{1,\dots,G+Z\}$, an affinity score is first computed as $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x},$ which is then normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias term introduced for auxiliary-loss-free load balancing (Liu et al., 2024a) (details in Appendix D.3.1). The router thus produces a set of active granularity indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}.$ That is, only the Top-$K$ granularity experts are activated, while null experts contribute no computation even when selected. By selectively activating granularity experts and allowing null experts to absorb routing probability mass, Kairos dynamically adapts the effective granularity resolution for each time series segment according to its local information, while maintaining a sparse and parameter-efficient architecture.
Step 2: Granularity-wise Encoding. For each selectively activated granularity expert at level $i$, we apply the corresponding patching and embedding operations. Specifically, a given time series segment $\mathbf{x}$ is re-partitioned into a sequence of patches at granularity $i$:
$\boldsymbol{p}_{i,1},\dots,\boldsymbol{p}_{i,M_{i}}]=\mathrm{Patchify}(\mathbf{x},i)\in\mathbb{R}^{M_{i}\times P_{i}},$ (1)
where $P_{i}$ denotes the patch size at granularity $i$, and $M_{i}=\left\lceil{P}/{P_{i}}\right\rceil$ is the resulting number of patches obtained via non-overlapping segmentation, with zero-padding applied when necessary. Each patch is then projected into the latent space to obtain an embedding sequence $\boldsymbol{E}_{i}=[\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,M_{i}}]$:
$\quad\boldsymbol{e}_{i,m}=\pi_{i}(\boldsymbol{p}_{i,m})\in\mathbb{R}^{D_{h}},m\in[1,M_{i}]\penalty 10000\ .$ (2)
where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer multilayer perceptron (MLP).
Step 3: Mixture-of-Size Fusion. In the final encoder stage, we fuse the embeddings produced by the activated granularity experts into a unified token sequence, serving as the input to the Heterogeneity-Aware Transformer. First, for a time-series segment $\mathbf{x}$, we align all granularity-specific embeddings $\boldsymbol{E}_{i}$ to the finest activated resolution $P_{\min}=\min\{P_{i}\mid i\in\mathcal{I}_{\mathrm{valid}}\},$ yielding a target sequence length $M=\lceil P/P_{\min}\rceil$. This alignment is crucial to eliminate representation bias arising from unequal sequence lengths induced by different patch sizes across granularities. To enable efficient alignment, we design the patch sizes to be nested, i.e., $P_{i+1}$ is divisible by $P_{i}$. Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix D.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i},\alpha_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
Sample-level Result. Finally, the representations from all coarse segments are concatenated to form the encoder output
$\boldsymbol{H}=\mathrm{Concat}(\bar{\boldsymbol{E}}_{1},\ldots,\bar{\boldsymbol{E}}_{N})\in\mathbb{R}^{\tilde{T}\times D_{h}},$ (3)
where $\tilde{T}=\sum_{n=1}^{N}M_{n}$ and $M_{n}$ is the embedding sequence length of the $n$-th segment, forming the representation sequence for the given time series sample $\mathbf{X}$.
Unlike prior approaches such as Chen et al. (2024); Zhang et al. (2024), which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level. (ii) decouples temporal heterogeneity from a fixed representational scale by explicitly and dynamically assigning patch size via routing and multi-granularity encoding, enabling more effective intra-patch representation learning. This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1. As a result, varying temporal patterns are modeled via the lightweight Mixture-of-Size Encoder, reducing the burden of subsequent model parameterization.
3.3 Heterogeneity-Aware Transformer
Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer (Vaswani et al., 2017) encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head. To inject temporal information into the self-attention mechanism, we propose Dynamic Rotary Position Embedding (DRoPE). Unlike static encodings such as Vaswani et al. (2017); Su et al. (2024), DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
Base Structure of DRoPE. Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) Su et al. (2024), where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers. Each pair is rotated by an angle proportional to the position $t$ and a fixed frequency $\theta_{d}$ as below, where $d\in[0,D_{h}/2-1]$ indexes the dimension:
$f_{\text{DRoPE}}(\boldsymbol{z},t)=(\boldsymbol{z}_{2d}+i\boldsymbol{z}_{2d+1})e^{it\theta_{d}}$ (4)
Here the frequencies are traditionally pre-defined as $\theta_{d}=b^{-2d/D_{h}}$. DRoPE generalizes this formulation by dynamically transforming both the frequency $\theta$ and the token position $t$. A detailed derivation of the RoPE calculation is provided in Appendix D.4.
Instance-Level Frequency Modulation. Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem D.1). We propose to modulate the base frequencies using instance-specific spectral features. Specifically, given a time series input $\mathbf{X}$, we extract the magnitudes of the first $\omega$ low-frequency components via Fast Fourier Transform, denoted as $\mathbf{X}_{\mathrm{fft}}$. A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$, to produce adaptive frequencies $\theta_{\mathrm{inst}}$:
$\theta_{\text{inst},d}=\gamma_{d}\cdot\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
This affine transformation allows Kairos to re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance.
Granularity-Aware Position Calibration. In our multi-granularity framework, the discrete token index $t$ is no longer an accurate proxy for time because different tokens represent different physical durations, a direct consequence of the granularity routing and tokenization in our Mixture-of-Size Encoder, as described in Section 3.2. To maintain temporal consistency across mixed patch sizes, we replace the integer index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
The resulting DRoPE operation unifies these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
$f_{\text{DRoPE}}(\boldsymbol{z},t)=(\boldsymbol{z}_{2d}+i\boldsymbol{z}_{2d+1})e^{i\cdot\tilde{t}\cdot\theta_{\text{inst},d}}$ (6)
By measuring relative distances in terms of absolute physical time and instance-specific frequencies, DRoPE enables the Transformer to learn consistent temporal dependencies considering the diverse encoding granularity and the intrinsic properties of the data sample.
3.4 Multi-Patch Decoder
To mitigate the error accumulation typical of patch-wise autoregressive decoding while maintaining structural flexibility, Kairos utilizes a Multi-Patch Decoder. This component forecasts $J$ future patches in parallel, effectively decoupling the prediction horizon from iterative dependencies. We first introduce a sequence of $J$ learnable forecast tokens, $\mathbf{Q}_{\text{forecast}}=\{\mathbf{q}_{1},\dots,\mathbf{q}_{J}\}$. Multi-Patch Decoder processes these token embeddings with the latent representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ from the above Heterogeneity-Aware Transformer through cross-attention, to produce a sequence of latent forecasting embeddings $\boldsymbol{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$.
@@ 3.4 Multi-Patch Decoder
Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block Das et al. (2024), to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
$\hat{\mathbf{Y}}_{(j-1)H_{s}+1:jH_{s}}=\text{ResidualFFN}(\boldsymbol{h}^{\text{de}}_{j}),\quad j\in[1,J]$ (7)
This multi-token design is strategically for two objectives regarding predominant forecasting paradigms. (i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps. (ii) Horizon flexibility: Unlike models that generate a single larger patch (Das et al., 2024), our approach offers superior granularity for variable prediction horizons. By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs. We provide a detailed analysis of this design in Appendix B. By combining the Mixture-of-Size Encoder for adaptive input processing with this Multi-Patch Decoder for flexible output generation, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.